OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is perpendicular to y = -4x -2 and passes through the point (-16, -11). Do not use spaces and enter fractions as improper reduced fractions when needed.

OpenStudy (anonymous):

Perpendicular lines have a slope of the negative reciprocal. If the slope of this line is 4, what is the slope of a perpendicular line?

OpenStudy (anonymous):

if a line is perpendicular then if you multiply the two slopes its suppose to give you -1. so -4x=-1 so x= 1/4 and x is the slope or gradient of the line we are looking for.

OpenStudy (anonymous):

I mean that if the slope is -4, what is the perpendicular slope?

OpenStudy (anonymous):

with point (-16,-11) and an arbitrary point (x,y) we look for the equation = (-16,-11) (x,y) and the slope 1/4 (y+16)/(x+11) = 1/4

OpenStudy (anonymous):

thus the equation will be?